OpenStudy (anonymous):

1. What is the simplest form of the radical expression? 4 fourth root of two x+ 6 fourth root of two x (1 point) 10fourth root of two x 10 fourth root of four x 20 fourth root of two x not possible to simplify

OpenStudy (anonymous):

\[\sqrt[4]{2x+6}\sqrt[4]{2x} = \sqrt[4]{2x(2x+6)}\]

OpenStudy (anonymous):

I havent really been taught this subject well could you explain? :) @monokerous

OpenStudy (anonymous):

\[10\sqrt[4]{2x}10\sqrt[4]{4x}20\sqrt[4]{2x}=10\times10\times20\sqrt[4]{2x \times4x \times2x}\]

OpenStudy (anonymous):

Basically what I did was times everything under the radical together, since they have the same root. After that, what you want to do is to simplify it.

OpenStudy (anonymous):

does that make sense? or would you like me to go further?

OpenStudy (anonymous):

No, thank you, I understand :) it makes perfect sense

OpenStudy (anonymous):

:) Great! You're welcome.